The state of stress at a point on a body is given by the following stress components: 0 = 15 MPa, Oy = -22 MPa and Try = 9 MPNote: the angle 0, defines the orientation of the maximum shear stress element with respect to the x- axis.
Answer #1

Gineni x= 15 MPa y = -22 MPa Txy 9 MPa Principal stressed 89,2= ² a toj + scogoodt + 4 Exy² لا - = = ( 15-22 + 1615 +22)2 +4x② Absoulute Tmax = larger q *5*112/19/13 1/17,073 = Larges -24,073 17-073 + 24.073 f 2 2 Absolute Maximum Shear stress a 20.5is shear stress Location of maximiem 14 ginen hen by tan 20 -1 (15 +22) 2 Exy 2x9 ofroy 0 = - 32.03 S 9 s*=htra 2 = 17.073 -2